OpenStudy (anonymous):

Help me PLEASE!! Question 1 (Essay Worth 10 points) (07.08) Part A: Divide (4x4y3 + 8x3y2 - 12x2y - 16x2y4) by -4x2y. Show your work, and justify each step. (6 points) Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? (2 points) Part C: What is the degree and classification of the polynomial you got in Part A? (2 points)

OpenStudy (anonymous):

@kirbykirby @SithsAndGiggles

OpenStudy (kirbykirby):

\[ \frac{4x^4y^3 + 8x^3y^2 - 12x^2y - 16x^2y^4}{-4x^2y}\] ?

OpenStudy (anonymous):

Hard xD,let me try..

OpenStudy (kirbykirby):

well I meant is this the correct expression :) ?

OpenStudy (anonymous):

ye

OpenStudy (kirbykirby):

Well I'll let you try as you said, and see what you get :P

OpenStudy (anonymous):

2x2y+8xy-3xy-4xy^4

OpenStudy (anonymous):

was hard for me,idk thats what I got :l

OpenStudy (aum):

It may be easier if you tery this: \(\Large \frac{a+b+c+d}{e} = \frac ae + \frac be + \frac ce + \frac de\)

OpenStudy (anonymous):

ugh

OpenStudy (kirbykirby):

I'm not sure exactly how you proceeded oye. Yes separating the fraction into 4 fractions might be easier to see how to simplify your expression. And don't forget the exponent rule: \[ \frac{x^a}{x^b}=x^{a-b}\] If you have two variables, it's the same idea: \[ \frac{x^ay^c}{x^by^d}=\frac{x^a}{x^b}\frac{y^c}{y^d}=x^{a-b}y^{c-d}\]

OpenStudy (aum):

\(\Large \frac{4x^4y^3 + 8x^3y^2 - 12x^2y - 16x^2y^4}{-4x^2y} = \) \(\Large \frac{4x^4y^3}{-4x^2y} + \frac{8x^3y^2}{-4x^2y} - \frac{12x^2y}{-4x^2y} - \frac{16x^2y^4}{-4x^2y} \) = ?

OpenStudy (anonymous):

keep dcing my bad

OpenStudy (anonymous):

40x^11y9 over -16x^8y

OpenStudy (kirbykirby):

is that a new question?

OpenStudy (anonymous):

no,the answer to his question xD

OpenStudy (anonymous):

im still trying to do part a...

OpenStudy (anonymous):

would this be the work for part a: (4x4y3 + 8x3y2 - 12x2y - 16x2y4) 4((x4y3 + 2x3y2 - 3x2y - 4x2y4)= 4 x^2 (x^2y^3 +2xy^2 -3y-4y^4)= 4 x^2 y (x^2y^2 +2xy -3-4y^3)= now divide by -4x^2y 4 x^2 y (x^2y^2 +2xy -3-4y^3) --------------------------- = - (x^2y^2 +2xy -3-4y^3) -4x^2 y

OpenStudy (anonymous):

wait uhm,I dont get what you want me to do..can you rewrite it like this but with the changes,idk if that is too much to ask sorry u__u

OpenStudy (kirbykirby):

ooh no never mind :o ...

OpenStudy (kirbykirby):

ignore what I said.. Its right lol

OpenStudy (anonymous):

thanks xDD

OpenStudy (kirbykirby):

well just distributing the negative: \(- (x^2y^2 +2xy -3-4y^3)=-x^2y^2-2xy+3+4y^3\)

OpenStudy (anonymous):

just to clarify which one is the answer? - (x^2y^2 +2xy -3-4y^3) -4x^2 y

OpenStudy (kirbykirby):

both \(- (x^2y^2 +2xy -3-4y^3)\) and \( -x^2y^2-2xy+3+4y^3\) are the same